Find the volume of a right circular cone obtained from a sector of a circle in which the radius is 26 cm and the central angle is 138.5 ° ? Round your final answer to 1 decimal place.
1
Expert's answer
2020-11-23T18:18:06-0500

"\displaystyle

\textsf{Slant height of cone} = \\
\textsf{Radius of circle which a sector is cut from.}\\

l = 26cm\\

\textsf{Circumference of base of cone} = \textsf{length of arc}\\

2\pi R = \frac{138.5\degree}{360\degree} \times 2\pi \times 26\\

R = \frac{3601}{360}\,cm\\

\begin{aligned}
V &= \frac{1}{3}\pi R^2 \sqrt{l^2 - R^2}
\\&= \frac{\pi}{3}\left(\frac{3601}{360}\right)^2 \times \sqrt{26^2 - \left(\frac{3601}{360}\right)^2}
\\&= 2514.5\,cm^3 \,(1 \,\textsf{d.p.})
\end{aligned}"